OpenStudy (anonymous):

The first and last terms of a 52-term arithmetic series are listed in the table. What is the sum of the series? Term Number Term 1 1 52 154

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Sum of an arithmetic series = N * (first term + last term) / 2

OpenStudy (ranga):

There are two formulas for the sum of an arithmetic series. In the previous problem we used the other formula. Here this formula is more appropriate.

OpenStudy (anonymous):

\[s _{n}=n/2(a1+a_n)\]

OpenStudy (ranga):

Yes. Identify each variable in the formula and plug in the numbers from the problem.

OpenStudy (anonymous):

a_1=1 a_n=52

OpenStudy (ranga):

Yes and what is n?

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

a_1=1 a_n=154 n = 52

OpenStudy (ranga):

Look at the left column in the problem. The first term number is 1. The last term number is 52. So how many terms are there?

OpenStudy (anonymous):

154

OpenStudy (ranga):

154 is the last term. But it is the fifty-second term. That means there are 52 terms in the series. n = 52.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i got 14300

OpenStudy (ranga):

Plug the numbers into the formula and calculate the sum.

OpenStudy (ranga):

Try again.

OpenStudy (anonymous):

3900

OpenStudy (ranga):

Are you just reading off your various answer choices?

OpenStudy (anonymous):

no im not

OpenStudy (anonymous):

am i still wrong

OpenStudy (ranga):

sum = 52 / 2 * (1 + 154) = ?

OpenStudy (anonymous):

52/2=26 1+154=155 26*155=4030

OpenStudy (anonymous):

i was doing it in my head i know where i made my mistake

OpenStudy (ranga):

yes, 4030.

OpenStudy (anonymous):

Thank you!

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

This was correct! Tysm!